User Tools

Site Tools


lbaops:lbajun2007:v210aholog

X-raid disks SWIN V001 BP1

Hobart will match the frequency/polarisation setup for DAS1 listed in the wiki table.

The log from experiment v210a has been copied to: ftp.atnf.csiro.au:/pub/people/vlbi/incoming/

ANTAB and UVFLG files have been copied to the same place.

There was a couple of wind stows and a power failure as recorded in the log comments.

Observers: Jamie Stevens then Simon Elingsen.

Observer comments follow:

2007.171.04:07:38.02;“weather: cool, 95% cloud cover, wind is 15km/h W

2007.171.11:40:08.74;”Coherence check

2007.171.13:34:17.24;“Wind stowed

2007.171.14:42:59.95;”Recording data again. Think we have recovered from power failure, doing system checks

2007.171.14:52:11.69;“Previous few tsys measurements spurious - a few attenuators needed reseting

2007.171.14:52:29.96;”after the power failure. The levels into the DAS were OK throughout though

2007.171.14:54:39.18;“Data from power failure resumption until now recorded with wrong DAS profile

2007.171.14:54:51.98;”(lost about 10 minutes of data). Profile now OK.

2007.171.14:57:02.36;“All systems appear to be OK now, have coherence etc.

2007.171.15:06:00.57;”wind stow

Regards,

Brett.

lbaops/lbajun2007/v210aholog.txt · Last modified: 2015/12/18 16:38 (external edit)